Two masses of 1 kg and 2 kg are moving with velocity 2m/s and 1m/s find the velocity of center of mass
1
Expert's answer
2020-09-21T08:28:29-0400

mass of particles are 1kg and 2 kg

Velocities of the particles are 2m/s and 1m/s


Velocity of center of mass is given by "V_{cm} = \frac{m_1v_1+m_2v_2}{m_1+m_2} = \frac{(1 * 2) + (2*1)}{1+2} = \frac{4}{3} m/s"